OpenStudy (anonymous):

Which is an example of the commutative property? A. 21 + (–21) = 0 B. 11 + (–3 + 8) = (11 + –3) + 8 C. –(12 + 7) = –12 + (–7) D. 13 + (–19) = –19 + 13

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

b + a = a + b and 13+(-19)=-19+13 looks the part.